OpenStudy (anonymous):

Please help me with algebra 2

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

Use composition of functions to determine if f(x) and g(x) are inverses. \[f(x) = \frac{ 3-x }{?x }\] \[g(x)=\frac{ 3 }{ x+1 }\]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

oops, ignore the question mark on the first function

OpenStudy (anonymous):

find \( f(g(x)) \)

OpenStudy (anonymous):

I know, but i'm not sure how to simplify f(g(x))

OpenStudy (anonymous):

then if u see \( f(g(x)) = x \) so they are inverse.

OpenStudy (anonymous):

if in fact they are inverses expect a whole orgy of cancellation leaving you with only \(x\)

OpenStudy (anonymous):

ok, \(\Large f(g(x)) = \frac{ 3-\frac{ 3 }{ x+1 } }{ x }\)

OpenStudy (anonymous):

ok,what have u done so far?

OpenStudy (anonymous):

I am having trouble simplifying, do I try to clear the fraction on the numerator???

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

do that

OpenStudy (anonymous):

get rid of the compound fraction by multiplying top and bottom by \(x+1\) would be my first choice there are other ways

OpenStudy (anonymous):

Sorry, i'm having trouble, if I multiply by x+1 how would the fractions clear?

OpenStudy (anonymous):

Can someone please show the work???

OpenStudy (anonymous):

\(\Large \frac{ 3-\frac{ 3 }{ x+1 } }{ x } = \frac{ \frac{ 3x+3-3}{ x+1 } }{ x } = \frac{ 3 }{ x+1 }\)

OpenStudy (anonymous):

which means they are NOT inverses maybe there was a typo somewheres

OpenStudy (anonymous):

Wait, shouldn't the denominator be 3/x+1

OpenStudy (anonymous):

yes it's 3\x+1

OpenStudy (anonymous):

so they're not inverse...

OpenStudy (anonymous):

No, you also have to plug in 3/x+1 to the denominator of f(x) i think

OpenStudy (anonymous):

no ;) i first pluged in g(x) :\( \Large f(g(x)) = \frac{ 3-\frac{ 3 }{ x+1 } }{ x } \)

OpenStudy (anonymous):

and found it that is equal to \( \Large \frac{ 3-\frac{ 3 }{ x+1 } }{ x } = \frac{ \frac{ 3x+3-3}{ x+1 } }{ x } = \frac{ 3 }{ x+1 } \)

OpenStudy (anonymous):

got it? @science00000

OpenStudy (anonymous):

I think you're wrong...

OpenStudy (anonymous):

why did you just leave the x on the denominator when you have to plug in g(x)

OpenStudy (anonymous):

\( \Large f(g(x)) = \frac{ 3-\frac{ 3 }{ x+1 } }{ x } \) did u get this that what i did?

OpenStudy (anonymous):

why did you only plug in g(x) to the numerator

OpenStudy (anonymous):

Oops ! specially thank u for ur awesome carefulness ;) yes

OpenStudy (anonymous):

in that way we can write : \(\Large f(g(x)) = \frac{ \frac{ 3x }{ x+1 } }{ \frac{ 3 }{ x+1 } }\)

OpenStudy (ajspeller):

Steps: 1) replace f(x) with the variable y. 2) change all of the x's to y's and the y's to x's 3) solve for y If the result is the other function, in this case g(x), then the function is an inverse.